The length is 2ft more than the width and is perimeter is 84 feet
Answer by checkley77(12844)

L=W+2
P=2L+2W
84=2(W+2)+2W
84=2W+4+2W
4W+4=84
4W=84-4
4W=80
W=80/4
W=20 ANS. FOR THE WIDTH.
L=20+2=22 ANS. FOR THE LENGTH.
PROOF:
84=2*22+2*20
84=44+40
84=84